# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@
Abstract
We introduce an algorithm that releases a pure differentially private sparse histogram over $n$ participants drawn from a domain of size $d\gg n$. Our method attains the optimal $\ell_{\infty}$-estimation error and runs in strictly $O(n\ln\ln d)$ time in the word-RAM model, thereby improving upon the previous best known deterministic-time bound of $\tilde{O}(n^{2})$ and resolving the open problem of breaking this quadratic barrier (Balcer and Vadhan, 2019). Central to our algorithm is a novel private item blanket technique with target-length padding, which transforms the approximate differentially private stability-based histogram algorithm into a pure differentially private one.
1 Introduction
@@
[…]
@@
Sampling Laplace noise requires representing real numbers on discrete machines, typically via double-precision floating-point arithmetic. As early as 2012, Mironov (2012) observed that due to finite precision and rounding artifacts, certain floating-point values cannot be generated. These artifacts can be exploited to distinguish exactly between neighboring inputs based on the outputs of the Laplace mechanism, thereby completely compromising its privacy guarantees. A fundamental mitigation is to avoid real-number arithmetic altogether by using discrete analogues of Laplace noise—such as the discrete Laplace distribution (Ghosh et al., 2009).
@@
Adding noise to all entries of $\vec{h}$ results in a running time proportional to $d$, which can be exponential in the input size. For example, consider the task of identifying popular URLs of up to 20 characters in length, [^1] Valid URL characters include digits (009999), letters (A–Z, a–z), and a few special characters (”-”, ”.”, ”_”, ”~”). as discussed in Fanti et al. (2016). This setting corresponds to a domain size of at least $d\geq 10^{36}$.
To avoid adding noise to all entries, a natural idea is to publish a sparse histogram $\tilde{h}$ to approximate the original histogram $\vec{h}$, since the latter is also sparse: it has at most $n$ nonzero entries. Some early works along this line, either satisfy only approximate differential privacy (Korolova et al., 2009; Bun et al., 2019a) —in contrast to the pure differential privacy guaranteed by the Laplace mechanism 222the distinction between pure and approximate differential privacy will be formally explained in Section 2 — or require sampling from complicated distributions (Cormode et al., 2012), for which it remains open whether one can sample efficiently while avoiding timing attacks (Balcer and Vadhan, 2019).
@@
Research Question: Can one reduce the nearly quadratic dependence on $n$ to nearly linear time, while preserving sparsity, privacy, and accuracy guarantees?
1.1 Our Contributions
Our contributions are twofold. First, we provide a positive answer to the open question posed by Balcer and Vadhan (2019). Second, as a byproduct, we develop a faster sampler for the relaxed discrete Laplace noise based on a general framework that may be of independent interest and applicable to other noise generation tasks. We discuss both results separately.
@@ 1.1 Our Contributions
Consistent with (Balcer and Vadhan, 2019), our construction builds on word-RAM model, and the privacy guarantee is respect to replacement neighboring relation, both formally defined in Section 2.
Theorem 1.1 (Private Sparse Histogram, Informal version of Theorem 4.1).
@@
Let $n,d,a_{\varepsilon},b_{\varepsilon}\in\mathbb{N}_{+}$ be integers that fit in a constant number of machine words, and define $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}\in\mathbb{Q}_{+}$. Given a histogram $\vec{h}\in[{0}\,.\,.\,{n}]^{d}$, there exists an $\varepsilon$-differentially private algorithm that runs deterministically in $O(n\ln\ln d)$ time and outputs a noisy histogram $\tilde{h}\in[{0}\,.\,.\,{n}]^{d}$ such that $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and  $\operatorname*{\mathbb{E}}\left[\lVert{\vec{h}-\tilde{h}}\rVert_{\infty}\right]\in O{({\frac{1}{\varepsilon}\cdot\ln d})}$.
The expected error bound above is asymptotically optimal, matching the lower bounds established in (Hardt and Talwar, 2010; Beimel et al., 2014; Balcer and Vadhan, 2019). A natural question is whether a similar guarantee can be achieved under the more stringent add/remove model of neighboring datasets for differential privacy. This setting is inherently more challenging than the replacement model—for example, the dataset size $n$ cannot be publicly revealed under the add/remove model, whereas it is permitted in the replacement model. Our answer is affirmative, albeit under a slightly relaxed notion: we design time-oblivious algorithms with expected linear, rather than worst-case linear, running time.
Our construction builds upon the recent framework of Ratliff and Vadhan (2025), which shows how to transform a pure DP, time-oblivious algorithm designed for inputs with a known upper bound on size into one that remains pure DP and time-oblivious without assuming any input size bound, under the add/remove neighboring model. We apply combine framework with our algorithm from Theorem 1.1, which serves as a key subroutine. The formal construction is presented in Section 4.3.
Applications. Our algorithm has a number of applications. For example, the sparse histogram problem is often formulated as a variant of the fundamental heavy hitter problem. Formally, a domain element $i\in[d]$ is said to be $\Delta$-heavy if $\vec{h}{\left[{i}\right]}\geq\Delta$. This variant seeks to identify all $\Delta$-heavy elements for $\Delta$ as small as possible (Bun et al., 2019b). Our algorithm is able to identify all $O{({\frac{1}{\varepsilon}\cdot\ln d})}$ heavy elements with high probability, as shown by the tail bound in Theorem 4.1. By the lower bound on $\ell_{\infty}$ estimation error, this is indeed the best one can hope to achieve.
Our algorithm also enables the first pure-DP sparse histogram algorithm with optimal $\ell_{\infty}$ error in the secure multiparty computation (MPC) model. We demonstrate a prototype protocol in Appendix D. In this setting, a group of participants (or servers) jointly compute a function without revealing their individual inputs, and the resulting computation satisfies differential privacy. This setting is often referred to as the distributed differential privacy (DDP) model. The shuffle model can be viewed as a special case of DDP model, in which the joint computation is restricted to randomly permuting locally perturbed messages. There is a long line of work on frequency estimation in the DDP model (Bittau et al., 2017; Balle et al., 2019; Cheu et al., 2019; Ghazi et al., 2021), culminating in a recent approximate-DP sparse histogram algorithm with optimal $\ell_{\infty}$ error (Bell et al., 2022). A key barrier to designing a pure-DP counterpart has been the lack of efficient, time-oblivious pure-DP algorithms (even in the central model), which are necessary to prevent timing side channels. Our algorithm overcomes this barrier.
@@
Our histogram protocol perturbs elements in the range $[{0}\,.\,.\,{n}]$ using a relaxed variant of discrete Laplace noise.
Theorem 1.2 (Relaxed Discrete Laplace, Informal version of Theorem 5.2).
Let $n,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma}\in\mathbb{N}_{+}$ be integers that fit in a constant number of machine words, such that $a_{\gamma}<b_{\gamma}$. Define $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}\in\mathbb{Q}_{+}$, $\gamma=a_{\gamma}/b_{\gamma}\in\mathbb{Q}_{+}\cap(0,1)$. There exists a randomized algorithm $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma})}}:\mathbb{N}_{\leq{n}}\rightarrow\mathbb{N}_{\leq{n}}$ with initialization time $\tilde{O}(\frac{1}{\varepsilon})$, memory usage $O\bigl{(}\frac{1}{\varepsilon}+\ln\frac{1}{\gamma}+\ln n\bigr{)}$, deterministic per-query running time $O(1)$ after initialization. It satisfies the following privacy and utility guarantees:
  • $\circ$
@@
The key building block of Theorem 1.2 is the following noise sampler.
Theorem 1.3 (Approximate Discrete Laplace Sampler, Informal Version of Theorem 5.3).
Let $a_{\varepsilon},b_{\varepsilon},a_{\delta},b_{\delta}\in\mathbb{N}_{+}$ be integers that fit in a constant number of machine words, where $a_{\delta}<b_{\delta}$, and define $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}\in\mathbb{Q}_{+}$, $\delta\doteq\frac{a_{\delta}}{b_{\delta}}\in\mathbb{Q}_{+}\cap{({0,1})}$. There exists a deterministic-time sampler with initialization time $\tilde{O}(1/\varepsilon)$, [^3] The $\tilde{O}$ notation hides dependence on $\ln(1/\varepsilon)$ and $\ln(1/\delta)$. memory usage $O\left(1/\varepsilon+\ln(1/\delta)\right)$, and $O(1)$ per-sample runtime, such that each sample is within total variation distance at most $\delta$ from a discrete Laplace noise $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$.
Directly applying the sampler does not yield a pure DP guarantee. This is not a major obstacle: following a technique introduced by Balcer and Vadhan (2019) (see Section 3), one can mix the sampler’s output distribution with a uniform one over $[{0}\,.\,.\,{n}]$ to convert an approximate DP guarantee into a pure one. Indeed, the sampler in Balcer and Vadhan (2019) adopts this exact strategy. Thus, the core challenge reduces to designing an efficient sampler whose output distribution is close in total variation distance to the discrete Laplace distribution.
Compared to prior deterministic-time samplers with the same sampling accuracy, our construction achieves the best known per-sample and batch running times, as summarized in Table 1. As we demonstrate in Section 5, our sampler is based on a general framework that instantiates the Alias method (Walker, 1977), with optimizations that leverage the special structure of the discrete Laplace distribution. This general framework is of independent interest and can be applied to other differentially private noise generation tasks.
Method Preprocessing Time Memory Time Per Sample Time for $m$ Samples
Theorem 5.3 $\tilde{O}{\big{(}{\frac{1}{\varepsilon}}\big{)}}$ $O{\big{(}{\frac{1}{\varepsilon}+\ln\frac{1}{\delta}}\big{)}}$ $O(1)$ $O(m)$
Table 1: Comparison of approximate discrete Laplace samplers within statistical distance $\delta$ of $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$, where $\varepsilon$ and $\delta$ are rational numbers whose reduced fractional representations fit within a constant number of machine words. The bounds for Canonne et al. (2020) are obtained by imposing a time limit on their sampler. The best running times are highlighted in orange. Per-sample running time may be of independent interest, especially when adapting samplers to scenarios where individual sampling time could leak information, such as in distributed settings.
Technical Overview
@@
Private Sparse Histogram. The algorithm has a remarkably simple structure with non-trivial analysis. It first adds a relaxed variant of discrete Laplace noise to the non-zero elements of $\vec{h}$, and selects those whose noisy counts exceed a specified threshold. For a neighboring histogram $\vec{h}^{\prime}$ with different support, this step might lead to a different set of selected elements.
@@
To hide this difference, the algorithm pads the selected set with uniformly random elements (without replacement) until its size reaches $n$. These random elements serve as a privacy blanket, and are drawn from all elements not selected in the first step—including both zero-count and possibly some non-zero elements. This sidesteps the costly procedure of identifying the zero elements in $\vec{h}$ with top-$n$ noisy counts, as required by Balcer and Vadhan (2019). Finally, for all selected elements, fresh noises are regenerated before releasing the output.
@@
As we detail in Section 4, the utility guarantee follows directly from the properties of the relaxed discrete Laplace noise, and the running time is inherited from our sampler’s efficiency (see Section 5). The privacy analysis, however, is the most delicate part. It requires proving that the distribution over the set of selected elements remains stable between neighboring histograms, which involves a careful case-by-case analysis.
@@
Noise Sampler. Our sampler construction (for Theorem 1.3) proceeds in two steps. First, we introduce a general framework for approximately sampling from arbitrary discrete distributions (not necessarily discrete Laplace) within a specified total variation distance. This framework builds on the classic Alias method (Walker, 1977), applied to a truncated support and fixed-point representations of probabilities. Despite its simplicity, it yields a deterministic-time sampler and improves upon the recent time-oblivious sampling scheme of Dov, David, Naor, and Tzalik (2023). The technique is generic and may extend to other noise distributions used in differential privacy, such as the discrete Gaussian.
@@
Second, we reduce sampling from the discrete Laplace distribution to sampling geometric noise. We exploit the fact that geometric noise has identical conditional distributions over intervals of equal length, allowing us to decompose it into two smaller geometric components. Each can be approximately sampled using our framework over reduced supports, thereby further lowering the space complexity.
Organization
@@
The remainder of the paper is organized as follows. Section 2 formally introduces the problem. Section 3 presents the relevant probability distributions and several building blocks used in our algorithms. Section 4 describes our private sparse histogram algorithm. We deliberately present this algorithm before introducing our noise sampling techniques in Section 5, to foreground the main result. This reordering does not affect the flow of the paper, as the histogram section is self-contained and can be read independently of the sampler’s implementation details. A protocol-style extension of the sparse histogram algorithm to the MPC setting appears in Appendix D. Finally, Section 6 provides a comprehensive review of related work.
2 Problem Description

### v2

# Optimal Pure Differentially Private Sparse Histograms in Deterministic Linear Time

@@
Abstract.
We present an algorithm that releases a pure differentially private (under the replacement neighboring relation) sparse histogram for $n$ participants over a domain of size $d\gg n$. Our method achieves the optimal $\ell_{\infty}$-estimation error and runs in strictly $O(n)$ time in the Word-RAM model, improving upon the previous best deterministic-time bound of $\tilde{O}(n^{2})$ and resolving the open problem of breaking this quadratic barrier (Balcer and Vadhan (2019)). Moreover, the algorithm admits an efficient circuit implementation, enabling the first near-linear communication and computation cost pure DP histogram MPC protocol with optimal $\ell_{\infty}$-estimation error. Central to our algorithm is a novel private item blanket technique with target-length padding, which hides differences in the supports of neighboring histograms while remaining efficiently implementable.
Differential Privacy, Sparse Histogram
doi: XXXXXXX.XXXXXXXccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Data anonymization and sanitizationccs: Theory of computation Design and analysis of algorithms
1. Introduction
@@
We focus on one of the most fundamental tasks in differential privacy: histogram publishing. In this setting, there are $n$ participants, each holding an element drawn from the domain [d]=[1..d][d]=[{1}\,.\,.\,{d}]. The histogram, $\vec{h}\in[{0}\,.\,.\,{n}]^{d}$, is a vector where $\vec{h}[i]$ equals the number of times element $i$ appears among the participants. The goal is to publish a differentially private version $\tilde{h}$ that well approximates $\vec{h}$ while protecting individual participant’s data.
@@
Sampling Laplace noise requires representing real numbers on discrete machines, typically via double-precision floating-point arithmetic. As early as (2012), Mironov (2012) observed that due to finite precision and rounding artifacts, certain floating-point values cannot be generated. These artifacts can be exploited to distinguish exactly between neighboring inputs based on the outputs of the Laplace mechanism, thereby completely compromising its privacy guarantees. A fundamental mitigation is to avoid real-number arithmetic altogether by using discrete analogues of Laplace noise—such as the discrete Laplace distribution (Ghosh et al., 2009).
@@
Adding noise to all entries of $\vec{h}$ results in a running time proportional to $d$, which can be exponential in the input size. For example, consider the task of identifying popular URLs of up to 20 characters in length, [^1] Valid URL characters include digits (099), letters (A–Z, a–z), and a few special characters (”-”, ”.”, ”_”, ”~”). as discussed in Fanti et al. (2016). This setting corresponds to a domain size of at least $d\geq 10^{36}$. To avoid adding noise to all entries, a natural idea is to publish a sparse histogram $\tilde{h}$ to approximate the original histogram $\vec{h}$, since the latter is also sparse: it has at most $n$ nonzero entries. Some early works along this line, either satisfy only approximate differential privacy (Korolova et al., 2009; Bun et al., 2019a)—in contrast to the pure differential privacy guaranteed by the Laplace mechanism  [^2] The distinction between pure and approximate differential privacy will be formally explained in Section 2—or require sampling from complicated binomial distributions (Cormode et al., 2012), for which it remains open whether one can sample efficiently while avoiding timing attacks (Balcer and Vadhan, 2019).
@@
Research Question: Can one design deterministic and linear-time pure differentially private histogram algorithms in the Word-RAM model while preserving privacy and accuracy guarantees?
1.1. Our Contributions
Our contributions are twofold. First, we provide a positive answer to the open question posed by Balcer and Vadhan (2019). Second, as a byproduct, we develop a general framework of deterministic-time approximate sampler for arbitrary discrete distributions, and instantiate it specifically for generating discrete Laplace noise. We discuss both results separately.
@@ 1.1. Our Contributions
Consistent with (Balcer and Vadhan, 2019), our construction builds on Word-RAM model, and the privacy guarantee is respect to replacement neighboring relation, both formally defined in Section˜2.
Algorithm $\ell_{\infty}$-Error Model Type Running Time
Laplace Mechanism (Dwork, 2006) $\Theta{({1/\varepsilon\cdot\ln d})}$ Real-RAM Deterministic $O(d)$
Geometric Mechanism (Ghosh et al., 2009) $\Theta{({1/\varepsilon\cdot\ln d})}$ Word-RAM Randomized Expected $O(d)$
Filter Mechanism (Cormode et al., 2012; Qiu and Yi, 2025) $\Theta{({1/\varepsilon\cdot\ln d})}$ Real-RAM Randomized Expected $O(n)$
PureSparseHistogram (Balcer and Vadhan, 2019) $\Theta{({1/\varepsilon\cdot\ln d})}$ Word-RAM Deterministic $\tilde{O}(n^{2})$
Theorem 1.1 $\Theta{({1/\varepsilon\cdot\ln d})}$ Word-RAM Deterministic $O(n)$
Table 1. Comparison of sparse histogram algorithms: error, model, type, and running time. The Real-RAM model assumes unit-cost arithmetic on infinite-precision floating-point numbers, while the Word-RAM model assumes unit-cost arithmetic on standard word-sized values.
Theorem 1.1 (Private Sparse Histogram, Informal version of Theorem˜4.1).
@@
Let $n,d\in\mathbb{N}_{+}$ and $\varepsilon\in\mathbb{Q}_{+}$, with representations that fit in a constant number of machine words. Given a dataset $\mathcal{X}$ of $n$ user-contributed elements from domain $[d]$, there exists an $\varepsilon$-differentially private algorithm that runs deterministically in $O(n)$ time and outputs a histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$ approximating the original histogram $\vec{h}$ of $\mathcal{X}$, such that $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and $\operatorname*{\mathbb{E}}\left[\lVert{\vec{h}-\tilde{h}}\rVert_{\infty}\right]\in O\!\left(\frac{1}{\varepsilon}\cdot\ln d\right)$.
The expected error bound above is asymptotically optimal, matching the lower bounds established in (Hardt and Talwar, 2010; Beimel et al., 2014; Balcer and Vadhan, 2019). Table 1 compares our algorithm wiht the previous ones. Beyond its applications to private data mining and machine learning, our algorithm also helps close a longstanding utility gap between the central and distributed differential privacy models for histograms.
From Central to Distributed DP: Matching Utility.
Our algorithm operates in the central DP model, where a trusted server collects participants’ data to release a private histogram. A commonly studied alternative is the distributed DP (DDP) model, where the server is untrusted. The simplest DDP setting is the local model, where each participant perturbs their own data before reporting it to the server. Unfortunately, this added privacy comes at a steep cost: local protocols incur substantial noise, and the best known frequency-estimation protocol achieves an expected $\ell_{\infty}$ error of $\Theta(\tfrac{1}{\varepsilon}\cdot\sqrt{n\ln d})$ (Bassily and Smith, 2015; Bun et al., 2019b; Wu and Wirth, 2022), creating a tight $\Omega(\sqrt{n})$ utility gap compared to the central model.
A significant body of work has attempted to bridge this gap by enabling participants (and, in some settings, multiple servers) to jointly and securely compute a function such that no party learns anything beyond their own input and the final DP output. Initial efforts focused on the shuffle model (Bittau et al., 2017; Balle et al., 2019; Cheu et al., 2019; Ghazi et al., 2021), where the joint computation is restricted to randomly permuting locally perturbed messages. Despite these advances, a gap remains: even the best near–linear-time shuffle protocols (Ghazi et al., 2021) still incur a polynomial (in $\log d$) degradation in utility compared to the optimal bounds achieved in the central model.
Recent works have further bridged this gap by exploiting the full potential of the DDP model, supporting richer classes of functionalities beyond shuffling via secure multiparty computation (MPC), as demonstrated by systems from Mozilla (Prio) (Corrigan-Gibbs and Boneh, 2017), Google (Bell et al., 2022), and Microsoft (Anderson et al., 2024). This line of work culminated in a recent $(\epsilon,\delta)$-DP sparse histogram algorithm (Bell et al., 2022) that achieves optimal $\ell_{\infty}$ error, on par with the best central-model guarantees. However, a Pure $\varepsilon$-DP counterpart with optimal error remains elusive. The primary obstacle has been the absence of efficient, time-oblivious Pure DP algorithms, even within the central model. Such algorithms are necessary in the MPC setting because data-dependent execution time or branching patterns can introduce side channels that compromise the privacy of the secret participant inputs. Our algorithm overcomes this barrier: it admits an efficient circuit implementation, enabling secure execution in the MPC setting using standard techniques (Ben-Or et al., 1988; Goldreich et al., 1987).
Theorem 1.2 (Private Sparse Histogram Circuit, Informal version of Theorem˜D.1).
Let $n,d\in\mathbb{N}_{+}$ and $\varepsilon\in\mathbb{Q}_{+}$, with representations that fit in a constant number of machine words. There exists a circuit with $\tilde{O}(n\omega)$ gates (where $\omega$ is the machine word size) implementing an $\varepsilon$-DP mechanism that takes a dataset $\mathcal{X}$ of $n$ user-contributed elements from domain $[d]$ as input, and outputs a histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$ approximating the original histogram $\vec{h}$ of $\mathcal{X}$, such that $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and $\operatorname*{\mathbb{E}}\left[\lVert{\vec{h}-\tilde{h}}\rVert_{\infty}\right]\in O{({\frac{1}{\varepsilon}\cdot\ln d})}$.
The number of gates scales with $n\omega$ rather than $n$ alone because, in the circuit model, merely reading $n$ input elements from $[d]$ (with each element described using $\log d\in O(\omega)$ bits) requires $O(n\omega)$ gates; hence, $n\omega$ is considered linear in the input size. Additionally, simulating Random Access Memory (RAM) operations of our algorithm in the circuit model introduces a logarithmic overhead in the gate complexity.
@@
We develop a general framework for constructing deterministic-time approximate samplers for arbitrary discrete distributions. This framework is of independent interest and applies broadly to differentially private noise-generation tasks. In addition, it strengthens the result of Dov, David, Naor, and Tzalik (2023) by achieving comparable running time and accuracy guarantees while using asymptotically less space.
Sampler Space Usage (words) Running Time / Sample TV Distance
[…]
Table 2. Comparison of deterministic approximate samplers for a discrete distribution $\mu$. Here, $\operatorname{C}_{\delta/2}$ denotes a smallest subset satisfying $\mu(\operatorname{C}_{\delta/2})\geq 1-\delta/2$, and $\omega$ is the word size. Our sampler improves space complexity by reducing its dependence on $\delta$ from $1/\delta$ to $\ln(1/\delta)$.
Theorem 1.3 (Deterministic Approximate Sampler, Informal Theorem˜B.2).
Let $\mu$ be a distribution over a discrete domain, let $\delta\in(0,1)$, and let $\operatorname{C}_{\delta/2}$ denote a smallest subset satisfying $\mu(\operatorname{C}_{\delta/2})\geq 1-\delta/2$. Then there exists a deterministic-time sampler for $\mu$ with the following guarantees:
  • $\triangleright$
[…]
  • $\triangleright$
    Sampling time per draw: $\frac{1}{\omega}\,{\big({\log\lvert{\operatorname{C}_{\delta/2}}\rvert+\log\tfrac{1}{\delta}}\big)}+O(1)$.
  • $\triangleright$
    Accuracy: The resulting distribution has total variation distance at most $\delta$ from $\mu$.
[…]
We now instantiate our general framework to generate a variant of discrete Laplace noise tailored for perturbing elements in the bounded range [0..n][{0}\,.\,.\,{n}]. This instantiation incorporates two key optimizations: (1) we adopt a technique from Balcer and Vadhan (2019) (see Section˜3) to mix the sampled noise with a uniform distribution over [0..n][{0}\,.\,.\,{n}], which ensures that the truncated output preserves pure differential privacy; (2) we exploit the memoryless property of the discrete Laplace distribution to further reduce the space required for sampling. The formal guarantees of this sampler are presented below and serve as a crucial building block for our linear-time private histogram protocol.
Theorem 1.4 (Relaxed Discrete Laplace, Informal version of Theorem˜B.3).
Let $n\in\mathbb{N}_{+}$, $\varepsilon\in\mathbb{Q}_{+}$ and $\gamma\in\mathbb{Q}_{+}\cap(0,1)$, with representations that fit in a constant number of machine words. There exists a randomized algorithm $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma})}}:\mathbb{N}_{\leq{n}}\rightarrow\mathbb{N}_{\leq{n}}$ with initialization time $\tilde{O}(\frac{1}{\varepsilon})$, memory usage $O\bigl(\frac{1}{\varepsilon}+\ln\frac{1}{\gamma}+\ln n\bigr)$, deterministic per-query running time $O(1)$ after initialization. It satisfies the following privacy and utility guarantees:
  • $\triangleright$
@@
Technical Overview.
@@
The algorithm has a remarkably simple structure with non-trivial analysis. It first adds a relaxed variant of discrete Laplace noise to the non-zero elements of $\vec{h}$, and selects those whose noisy counts exceed a specified threshold. For a neighboring histogram $\vec{h}^{\prime}$ with different support, this step might lead to a different set of selected elements.
@@
To hide this difference, the algorithm pads the selected set with uniformly random elements (without replacement) until its size reaches $n$. These random elements serve as a privacy blanket, and are drawn from all elements not selected in the first step—including both zero-count and possibly some non-zero elements. This sidesteps costly procedures such as the binomial-sampling step of Cormode et al. (2012), and the identification of top-$n$ noisy counts among zero elements in $\vec{h}$, as used by Balcer and Vadhan (2019). Finally, for all selected elements, fresh noises are regenerated before releasing the output.
@@
As we detail in Section˜4, the utility guarantee follows directly from the properties of the relaxed discrete Laplace noise, and the running time is inherited from our sampler’s efficiency (Theorem 1.4). The privacy analysis, however, is the most delicate part. It requires proving that the distribution over the set of selected elements remains stable between neighboring histograms, which involves a careful case-by-case analysis.
Noise Sampler.
@@
Our general framework for approximately sampling from arbitrary discrete distributions builds directly on the classic Alias method (Walker, 1977), applied to carefully chosen truncated supports and truncated probability representations to achieve a prescribed total variation distance. Despite its simplicity, this yields a deterministic-time sampler and improves upon the recent time-oblivious sampling scheme of Dov, David, Naor, and Tzalik (2023), making it worthwhile to document in this context.
@@
Next, we instantiate the framework for the discrete Laplace distribution. Specifically, we reduce sampling from the discrete Laplace to sampling geometric noise, exploiting the fact that the discrete Laplace is almost a two-sided geometric distribution. We further leverage the property that geometric noise has identical conditional distributions over intervals of equal length, allowing decomposition into two smaller geometric components. Finally, following a technique by Balcer and Vadhan (2019) (see Section˜3), we mix the sampled noise with a uniform distribution over [0..n][{0}\,.\,.\,{n}] to ensure that the truncated output preserves pure differential privacy.
Organization.
@@
The remainder of the paper is organized as follows. Section˜2 formally introduces the problem. Section˜3 reviews the relevant probability distributions and a key building block used in our algorithms. Due to space constraints, we present only our private sparse histogram algorithm in Section˜4 and defer the description of our noise-sampling procedures to Appendix˜B. This reordering does not affect the exposition, as Section˜4 is self-contained and can be read independently of the sampler implementation details. The circuit realization of our sparse histogram algorithm appears in Appendix˜D. Finally, Appendix˜F provides a comprehensive overview of related work.
2. Problem Description